OpenStudy (anonymous):

the perimeter of a rectangle carpet is 70 ft. the width is three-fourths the length. find the width

OpenStudy (anonymous):

\[W = \frac{3}{4}L \implies L = \frac{4}{3}W\] Now use: \[Perimeter = 2(L + W)\] \[70 = 2 (\frac{4}{3}W + W)\] Solve it for W..

OpenStudy (anonymous):

\[\frac{4}{3}W + W = \frac{4 + 3}{3}W \implies \frac{7}{3}W\] So you are left with: \[70 = 2 \times \frac{7}{3}W\] Now can you solve it??

OpenStudy (anonymous):

is the answer 12.35

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

There in the last equation, multiply by 3 both the sides first and tell me what are you getting??

OpenStudy (anonymous):

\[W = \frac{70 \times 3}{7 \times 2}\] Can you solve it??

OpenStudy (anonymous):

is it 15

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome dear..